# Bernstein-type Inequalities for Markov Chains and Markov Processes: A Simple and Robust Proof

### v3

@@ Abstract.
We establish a new Bernstein-type deviation inequality for general discrete-time Markov chains via an elementary approach. More robust than existing works in the literature, our result only requires the Markov chain to satisfy an iterated Poincaré inequality. Moreover, our method can be readily generalized to continuous-time Markov processes.
@@ 1. Introduction
Concentration inequalities describe the probability that the average of a sequence of random variables is close to its expected value. Motived by the profound impact of these inequalities on many scientific applications, significant efforts have been made to extend concentration results from independent random variables to dependent ones. In particular, the concentration of Markov chains has been an extensively studied topic. Let $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ be a Markov chain on state space $\Omega$ with a unique invariant distribution $\mu$. Let $L_{2,\mu}$ be the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product $\langle f,g\rangle_{\mu}=\mathbb{E}_{\mu}[f\cdot g]$, and let $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ be the mean-zero subspace of $L_{2,\mu}$. Our goal of this paper is to bound the deviation probability $\mathbb{P}\left\{{|\frac{1}{n}\sum_{k=1}^{n}f(Z_{k})|\geq\delta}\right\}$ for any bounded function $f\in L_{2,\mu}^{0}$.
In general, the convergence speed of the empirical mean $\frac{1}{n}\sum_{k=1}^{n}f(Z_{k})$ depends on certain spectral properties of the corresponding transition operator $\bm{P}$. In this paper, we use the smallest nonzero singular value of $\bm{\mathbf{I}}-\bm{P}$ to quantify the ergodicity of $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ and establish a Bernstein-type inequality for Markov chains with a nonzero iterated Poincaré gap. To be specific, we define the iterated Poincaré gap (IP gap) of $\bm{P}$ as
@@
Then for any $\delta>0$ and any function $f$ such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$, we can obtain the following Bernstein-type tail bound:
@@
There have been numerous studies that utilize certain spectral gaps of the transition operators to derive concentration inequalities for Markov chains [5, 6, 2, 7, 10, 8, 3, 4]. When $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ is reversible, the corresponding $\bm{P}$ is self-adjoint under the $L_{2,\mu}$ inner product. In this case, the spectral gap of $\bm{P}$ is given by
@@
When the number of states $|\Omega|$ is finite, $\eta$ is just the gap between 1 and the second largest eigenvalue of $\bm{P}$. It is well-known that, in the reversible case, a nonzero spectral gap of $\bm{P}$ implies ergodicity.
For nonreversible Markov chains, various notions of spectral gaps have been proposed to generalize the ordinary spectral gap. A commonly used one is the absolute spectral gap $\eta_{a}$, which is defined as
@@
To the best of our knowledge, most of the existing concentration inequalities for nonreversible Markov chains require a nonzero $\eta_{a}$. Unfortunately, the existence of nonzero absolute gaps cannot be guaranteed even for irreducible Markov chains (A simple example is shown in Appendix A). A variant of the absolute spectral gap, called the pseudo spectral gap, was first introduced in [8] to establish concentration inequalities. More precisely, let $\bm{P}^{*}$ be the adjoint of $\bm{P}$ under the $L_{2,\mu}$ inner product, then the pseudo spectral gap of $\bm{P}$ is defined as the supremum, over all $k\in\mathbb{N}_{+}$, of the spectral gap of $(\bm{P}^{*})^{k}\bm{P}^{k}/k$. This quantity, however, could be difficult to compute in practice. Advanced theoretical studies, estimations, and applications of the pseudo spectral gap can be found in [11, 12].
Another way to define the spectral gap for a nonreversible $\bm{P}$ is by considering the additive symmetrization of $\bm{P}$. To be specific, the symmetric spectral gap of $\bm{P}$ is defined as
@@
which can be seen as a natural generalization of (1.4). This quantity is employed to develop concentration inequalities in [3, 4]. However, there seems to be a technical gap in their proof (which will be explained and discussed in Appendix B).
The reason we choose to use the IP gap is that the relations $\eta_{p}\geq\eta_{s}\geq\eta_{a}$ always hold for any Markov transition operator $\bm{P}$. Furthermore, for any finite-state irreducible Markov chain, the corresponding IP gap $\eta_{p}$ is always nonzero and thus is more robust for quantifying ergodicity. Consequently, our result is more user-friendly and applicable to a mush wider range of Markov chains. In fact, the IP gap was first introduced in [1] to study the ergodicity of nonreversible Markov chains. We refer the reader to [1] for a more comprehensive introduction.
We establish (1.2) by recursively bounding the moment-generating function of $\sum_{k=1}^{n}f(Z_{k})$ in a way that only an iterated Poincaré inequality is needed, providing a more straightforward and elementary method compared to existing approaches. Furthermore, our method can be naturally generalized to continuous-time Markov processes. To be specific, let $(Z_{t})_{t\geq 0}$ be a Markov process on $\Omega$ with an invariant distribution $\mu$, and let $f$ be a function that satisfies the same assumptions as above. In this case, we can similarly define the IP gap $\eta_{p}$ for the associated Markov semigroup via an iterated Poincaré inequality and establish the following Bernstein-type inequality:
@@ 2.1. Setting
Throughout this paper, $\Omega$ is a Polish space equipped with a probability measure $\mu$. $L_{2,\mu}$ represents the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product
@@
and $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ represents the mean-zero subspace of $L_{2,\mu}$.
@@ 2.2. Exponential moment bound
We will first consider the case $\theta=1$. After that, we can bound $m(\theta)$ for arbitrary $\theta>0$ by replacing the function $f$ with $\theta f$. For $n\geq 1$, define
@@
$\bm{L}G_{n}(z)$ $=\bm{P}G_{n}(z)-G_{n}(z)$
$=\operatorname{\mathbb{E}}\left[G_{n}(Z_{2})\,|\,Z_{1}=z\right]-G_{n}(z)$
$=\operatorname{\mathbb{E}}\big{[}\mathrm{e}^{F(Z_{2:(n+1)})}-\mathrm{e}^{F(Z_{1:n})}\,\big{|}\,Z_{1}=z\big{]}$
$=\operatorname{\mathbb{E}}\big{[}(f(Z_{n+1})-f(Z_{1}))\int_{0}^{1}\mathrm{e}^{sF(Z_{2:(n+1)})+(1-s)F(Z_{1:n})}\,\mathrm{d}{s}\,\big{|}\,Z_{1}=z\big{]},$
@@
With the above lemma, we can bound $a_{n}$ as follows when $f$ is sufficiently small.
@@
Let $f:\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Suppose that $01-4M^{2}/\eta_{p}^{2}>$ 0, then for $n\in\mathbb{N}_{+}$,
@@
By the definition of $\eta_{p}$, it is apparent that $\eta_{p}\leq 2$. Owing to the assumption $01-4M^{2}/\eta_{p}^{2}>$ 0, one has $M<\eta_{p}/2\leq 1$. We shall bound $a_{n}$ by recursion. For any $n\geq 2$, we have
@@
$\big{(}\mathrm{e}^{2M}-\mathrm{e}^{M}\big{)}/M-1\leq 4M,\quad\text{for any $M\in(0,1]$}.$ (2.10)
@@
To continue, we use the above to compute that
@@
Invoking the assumption $c^{2}=1-4M^{2}/\eta_{p}^{2}>0$, we obtain $\operatorname{\mathbb{E}}_{\mu}[G_{n}^{2}]\leq a_{n}^{2}/c^{2}$. Finally, we have
$\left(\operatorname{\mathbb{E}}_{\mu}\big{[}(\bm{L}G_{n})^{2}\big{]}\right)^{1/2}\leq 2M\left(\operatorname{\mathbb{E}}_{\mu}\big{[}G_{n}^{2}\big{]}\right)^{1/2}\leq\frac{2M}{c}\cdot a_{n}.$
We then combine (2.9), (2.11), and (2.12) to obtain
@@
We have used the fact that $\sqrt{1-c^{2}}+c\leq\sqrt{2}<3/2$ for any $c\in[0,1]$. Unrolling the recursion yields $a_{n}\leq\mathrm{e}^{(n-1)\sigma M(2+6\eta_{p})/(c\eta_{p})}a_{1}$. In fact, using the same argument we can show that $a_{1}\leq\mathrm{e}^{\sigma M(2+6\eta_{p})/(c\eta_{p})}$. Therefore, we have $a_{n}\leq\mathrm{e}^{n\sigma M(2+6\eta_{p})/(c\eta_{p})}$ for all $n\geq 1$. ∎

### v4

@@ Abstract.
We establish a new Bernstein-type deviation inequality for general (non-reversible) discrete-time Markov chains via an elementary approach. More robust than existing works in the literature, our result only requires the Markov chain to satisfy an iterated Poincaré inequality. Moreover, our method can be readily generalized to continuous-time Markov processes.
@@ 1. Introduction
Concentration inequalities describe the probability that the average of a sequence of random variables is close to its expected value. Motivated by the profound impact of these inequalities on many scientific applications, significant efforts have been made to extend concentration results from independent random variables to dependent ones. In particular, the concentration of Markov chains has been an extensively studied topic. Let $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ be a discrete-time Markov chain on a state space $\Omega$ with a unique invariant distribution $\mu$. Let $L_{2,\mu}$ be the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product $\langle f,g\rangle_{\mu}=\mathbb{E}_{\mu}[f\cdot g]$, and let $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ be the mean-zero subspace of $L_{2,\mu}$. Our goal of this paper is to bound the deviation probability $\mathbb{P}\left\{{|\frac{1}{n}\sum_{k=1}^{n}f(Z_{k})|\geq\delta}\right\}$ for any bounded function $f\in L_{2,\mu}^{0}$.
In general, the convergence speed of the empirical mean $\frac{1}{n}\sum_{k=1}^{n}f(Z_{k})$ depends on certain spectral properties of the corresponding transition operator $\bm{P}$. In particular, ergodicity (in different senses) for non-reversible Markov chains has been proved to be guaranteed and quantified by various notions of spectral gaps of $\bm{P}$ [3, 12, 7, 13, 1]. In this paper, we use the smallest nonzero singular value of $\bm{\mathbf{I}}-\bm{P}$ (with respect to $L_{2,\mu}$) to quantify the ergodicity of $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ and establish a Bernstein-type inequality for Markov chains with a nonzero iterated Poincaré gap. To be specific, we define the iterated Poincaré gap (IP gap) of $\bm{P}$ as
@@
Here and below, $\|h\|_{\mu}=\langle h,h\rangle_{\mu}^{1/2}$. Then for any $\delta>0$ and any function $f$ such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$, we can obtain the following Bernstein-type tail bound:
@@
The definition of this IP gap coincides with that of the non-reversible spectral gap first formally introduced by Chatterjee in [1], where he used this quantity to establish ergodicity for non-reversible Markov chains, generalizing similar results for reversible Markov chains based on the usual spectral gap. For instance, Chatterjee showed that the empirical variance along a non-reversible Markov chain of length $n$ can be bounded by $1/(n\eta_{p})$, in the same spirit as in the reversible case. The reader may thus also refer to the IP gap as the Chatterjee gap or the non-reversible spectral gap and can find a more comprehensive introduction to the topic in [1]. Nevertheless, we want to remark that the proof ideas and the results of this paper are novel and were developed independently and unaware of the work [1] by Chatterjee [^1] The core techniques and the main results of this paper were already established by the first author DH in 2020 and were introduced to Joel Tropp (Caltech) in a private communication in the same year. We then spent a few years trying to improve the estimates and to extend the results to non-commutative (matrix) settings (though failed) before we finally released this paper in 2024., and we use the name “iterated Poincaré gap” to emphasize the critical but simple use of the iterated Poincaré inequality in our proof.
There have been numerous studies that establish concentration inequalities for Markov chains in terms of certain spectral gaps of the transition operators [8, 9, 2, 11, 13, 15, 6, 4]. When $\{Z_{k}\}_{k\in\mathbb{N}_{+}}$ is reversible, the corresponding $\bm{P}$ is self-adjoint under the $L_{2,\mu}$ inner product. In this case, the spectral gap of $\bm{P}$ is given by
@@
In other words, $\eta$ is the gap between $1$ and the second largest eigenvalue of $\bm{P}$ (when the spectrum of $\bm{P}$ is well-defined). It is well-known that, in the reversible case, a nonzero spectral gap of $\bm{P}$ implies ergodicity and concentration of the Markov chain.
For non-reversible Markov chains, various notions of spectral gaps have been proposed as substitutes for the usual spectral gap in the study of concentration inequalities. A commonly used one is the absolute $L_{2}$-spectral gap (or simply the absolute spectral gap), which is defined as
@@
If $\bm{P}$ is considered as a linear operator on $L_{2,\mu}$, then $\eta_{a}$ is the gap between $1$ and the operator norm of $\bm{P}$ over the subspace $L_{2,\mu}^{0}$ (the latter is also the second largest singular value of $\bm{P}$ under the $L_{2,\mu}$ inner product). Alternatively, $\eta_{a}$ can be understood as the usual spectral gap of the self-adjoint operator $(\bm{P}^{*}\bm{P})^{1/2}$, where $\bm{P}^{*}$ is the adjoint of $\bm{P}$ with respect to the $L_{2,\mu}$ inner product. Moreover, in some literature (e.g. [8]) an alternative definition,
$\eta_{m}:=1-\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}h\|_{\mu}^{2}}{\|h\|_{\mu}^{2}},$ (1.6)
is employed, which is the usual spectral gap of $\bm{P}^{*}\bm{P}$, the multiplicative reversiblization of $\bm{P}$ (see also [3, 12, 1]). It is easily seen that these two gaps $\eta_{a}$ and $\eta_{m}$ are always comparable since $\eta_{m}=1-(1-\eta_{a})^{2}=(2-\eta_{a})\eta_{a}$ by definition, so it does not matter which definition is used. To the best of our knowledge, most of the existing concentration inequalities for non-reversible Markov chains require a nonzero $\eta_{a}$ (or $\eta_{m}$) [8, 11, 15, 6, 4]. Unfortunately, though a nonzero $\eta_{a}$ implies ergodicity of the Markov chain, the reverse is not true in general: a nonzero absolute spectral gap cannot be guaranteed even for irreducible Markov chains (Simple examples can be found in Appendix A).
Note that in some literature [13, 10, 1] the name “absolute spectral gap” refers to a different quantity defined as
$\gamma_{a}:=1-|\tilde{\lambda}_{2}(\bm{P})|,$ (1.7)
where $\tilde{\lambda}_{2}(\bm{P})$ denotes the second largest eigenvalue of $\bm{P}$ in absolute value. This definition of $\gamma_{a}$ coincides with that of $\eta_{a}$ in (1.5) for reversible Markov chains. However, it is not so favorable to establish concentration inequalities in terms of $\gamma_{a}$ for non-reversible Markov chains. In principle, one can first bound the deviation probability (the left-hand side of (1.2)) by $\exp(-\delta^{2}/O(t_{\mathrm{mix}}))$ (see e.g. [13]), where $t_{\mathrm{mix}}$ is the so called mixing time of the Markov chain, and then bound $t_{\mathrm{mix}}$ from above by a multiple of $1/\gamma_{a}$ [13, 5]. Nevertheless, it is proved in [5] that $t_{\mathrm{mix}}\lesssim|\Omega|/\gamma_{a}$ (where $|\Omega|$ is the size of the state space $\Omega$) for general Markov chains with an example showing the sharpness of this bound. Such undesirable dependence on the size of the state space makes $\gamma_{a}$ a less useful quantity in the non-reversible case. Therefore, we will not discuss $\gamma_{a}$ in this paper, and we will always refer to $\eta_{a}$ defined in (1.5) as the absolute spectral gap.
A generalization of the absolute spectral gap, called the pseudo spectral gap, was first introduced in [13] to establish concentration inequalities for non-reversible Markov chains. It is defined as
$\eta_{ps}:=\sup_{k\in\mathbb{N}_{+}}\frac{1}{k}\left(1-\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}^{k}h\|_{\mu}^{2}}{\|h\|_{\mu}^{2}}\right)=\sup_{k\in\mathbb{N}_{+}}\frac{\eta\big((\bm{P}^{*})^{k}\bm{P}^{k}\big)}{k},$ (1.8)
where $\eta((\bm{P}^{*})^{k}\bm{P}^{k})$ is the usual spectral gap of the self-adjoint operator $(\bm{P}^{*})^{k}\bm{P}^{k}$. Though it could be difficult to compute this quantity exactly, one can obtain an easy lower bound by taking the supremum over $k\in\{1,\dots,K\}$ for some finite $K$. Nevertheless, this relatively more complicated definition of the pseudo spectral gap makes it less user-friendly in practice. Advanced theoretical studies, estimations, and applications of the pseudo spectral gap can be found in [16, 17].
Another way to define the spectral gap for a non-reversible Markov chain is by simply using the usual spectral gap of the self-adjoint operator $(\bm{P}+\bm{P}^{*})/2$, the additive symmetrization of $\bm{P}$ (see [3, 12, 1]). To be specific, the symmetric spectral gap of $\bm{P}$ is defined as
@@
which can be seen as a natural generalization of (1.4). This quantity is employed to obtain concentration inequalities in [4, 6]. However, there seems to be a technical gap in their proof of a critical lemma involving the symmetric spectral gap (which will be explained and discussed in Appendix B). Other than these unjustified results, the symmetric spectral gap has not been used in concentration inequalities for non-reversible Markov chains.
Note that in the reversible case, it always holds that $\eta=\eta_{p}=\eta_{s}$, and if the second largest eigenvalue of $\bm{P}$ in absolute value is positive, then it also holds that $\eta_{p}=\eta_{a}$. The reason we prefer the IP gap is that the relations $\eta_{p}\geq\eta_{s}\geq\eta_{a}$ (Lemma A.1) and $\eta_{p}\geq\eta_{ps}/2$ (see [1]) always hold for any non-reversible Markov chain, meaning that the IP gap is the most robust one among all these commonly used gaps for quantifying ergodicity. In fact, for any finite-state irreducible Markov chain, the corresponding IP gap $\eta_{p}$ is always nonzero. Moreover, there are simple examples where $\eta_{p}>0$ but the ratios $\eta_{a}/\eta_{p}$ and $\eta_{ps}/\eta_{p}$ can be arbitrarily small or even zero (see Appendix A). Consequently, our results are more robust, user-friendly, and applicable to a much wider range of Markov chains.
We establish (1.2) by recursively bounding the moment-generating function of $\sum_{k=1}^{n}f(Z_{k})$ in a way that only an iterated Poincaré inequality is needed, providing a more straightforward and elementary method compared to existing approaches. Furthermore, our method can be naturally generalized to continuous-time Markov processes. To be specific, let $(Z_{t})_{t\geq 0}$ be a continuous-time Markov process on $\Omega$ with an invariant distribution $\mu$, and let $(\bm{P}_{t})_{t\geq 0}$ be the associated Markov semigroup. In this case, we can similarly define the IP gap $\eta_{p}$ for $(\bm{P}_{t})_{t\geq 0}$ via an iterated Poincaré inequality and establish the following Bernstein-type inequality for any function $f$ that satisfies the same assumptions as above:
@@
We remark that though the proof in the continuous-time case is parallel to that in the discrete-time case, the former is cleaner and easier to read for some technical reasons. We hence suggest reading the proof in the continuous-time case first to quickly get to the key idea of our approach.
@@ 2.1. Setting
Throughout this paper, $\Omega$ is a Polish space equipped with a probability measure $\mu$. $L_{2,\mu}$ denotes the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product
@@
and $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ denotes the mean-zero subspace of $L_{2,\mu}$. We define $\|f\|_{\mu}:=\langle f,f\rangle_{\mu}^{1/2}$. Correspondingly, we have
$\operatorname{Var}_{\mu}[f]=\operatorname{\mathbb{E}}_{\mu}[(f-\operatorname{\mathbb{E}}_{\mu}f)^{2}]=\|f-\operatorname{\mathbb{E}}_{\mu}f\|_{\mu}^{2}.$
@@ 2.2. Exponential moment bound
To simplify notations, we will first consider the case $\theta=1$. After that, we can bound $m(\theta)$ for arbitrary $\theta>0$ by replacing the function $f$ with $\theta f$. For $n\geq 1$, define
@@
$\bm{L}G_{n}(z)$ $=\operatorname{\mathbb{E}}\left[G_{n}(Z_{2})\,|\,Z_{1}=z\right]-G_{n}(z)$
$=\operatorname{\mathbb{E}}\left[\mathrm{e}^{F(Z_{2:(n+1)})}-\mathrm{e}^{F(Z_{1:n})}\,\big|\,Z_{1}=z\right]$
$=\operatorname{\mathbb{E}}\left[(f(Z_{n+1})-f(Z_{1}))\int_{0}^{1}\mathrm{e}^{sF(Z_{2:(n+1)})+(1-s)F(Z_{1:n})}\,\mathrm{d}{s}\,\Big|\,Z_{1}=z\right],$
@@
With the above lemma, we can bound $a_{n}$ as follows given that $f$ satisfies some smallness condition.
@@
Let $f:\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Suppose that $1-4M^{2}/\eta_{p}^{2}>0$, then for any $n\in\mathbb{N}_{+}$,
@@
By the definition of $\eta_{p}$, it is apparent that $\eta_{p}\leq 2$. Owing to the assumption $1-4M^{2}/\eta_{p}^{2}>0$, one has $M<\eta_{p}/2\leq 1$. We shall bound $a_{n}$ by recursion. For any $n\geq 1$, we have
@@
$\frac{\mathrm{e}^{2M}-\mathrm{e}^{M}}{M}-1\leq 4M,\quad\text{for any $M\in(0,1]$}.$ (2.10)
@@
To continue, we use the above result to compute that
@@
Using the assumption $c^{2}=1-4M^{2}/\eta_{p}^{2}>0$, we obtain
$\operatorname{\mathbb{E}}_{\mu}[G_{n}^{2}]\leq\frac{a_{n}^{2}}{c^{2}}.$
Finally, we have
$\left(\operatorname{\mathbb{E}}_{\mu}\big[(\bm{L}G_{n})^{2}\big]\right)^{1/2}\leq 2M\left(\operatorname{\mathbb{E}}_{\mu}\big[G_{n}^{2}\big]\right)^{1/2}\leq\frac{2M}{c}\cdot a_{n}.$ (2.13)
We then combine (2.9), (2.11), (2.12), and the above to obtain
@@
We have used the fact that $\sqrt{1-c^{2}}+c\leq\sqrt{2}<3/2$ for any $c\in[0,1]$. Unrolling this recursion yields $a_{n}\leq\mathrm{e}^{(n-1)\sigma M(2+6\eta_{p})/(c\eta_{p})}a_{1}$. In fact, using the same argument we can show that $a_{1}\leq\mathrm{e}^{\sigma M(2+6\eta_{p})/(c\eta_{p})}$. Finally, we obtain
$a_{n}\leq\exp\left(n\sigma M\cdot\frac{2+6\eta_{p}}{c\,\eta_{p}}\right)$
for all $n\geq 1$. ∎